A 12-V battery causes a current of 0.80A  through a resistor.

A) What is its resistance?

B) How many joules of energy does the battery lose in a minute?


You buy a 75-W lightbulb in Europe, where electricity is delivered to homes at 240 V.


If you use the lightbulb in the United States at 120 V (assume its resistance does not change), how bright will it be relative to 75-W 120-V bulbs? [Hint: Assume roughly that brightness is proportional to power consumed.]         
Answer #1

A) R= V/I = 12/.8 = 15 ohm


B) Energy = power * time = V*I*60 = 12*.8*60= 576 j


if voltage is halved in US, the power consumed would be 1/4 times the orignal, as P =V^2/R, therefore, the brightness would be one fourth
Answer #2

p=ie, or watts=volts times amps
12 x .8 = 9.6 watts
A) r = 12/0.8 = 15 ohms (V = I.R)

B) energy = 9.6w x 60 min = 576 joules



U = RI, and , P = RI